Mercury is added to a cylindrical container to a d
Answer #1

prerssure at the bottom of the water is P2 = P1+(rho_w*g*(0.7-d))

P2 = 1.013*10^5 +(1000*9.8*(0.7-d))

pressure at the bottom of the mercury is P

P = P2 + (rho_hg*g*d)

P = 1.013*10^5 +(1000*9.8*(0.7-d)) + (rho_hg*g*d)

1.2*1.013*10^5 = 1.013*10^5 +(1000*9.8*(0.7-d)) + (1.36*10^4*9.8*d)


0.2*1.013*10^5 = (6860 - 9800*d + 133280*d)


d = 0.108 m